Back to the NIKA2 main page

Daily Reports

Jan. 17th, Tuesday

Frederic, Sofia, Angelos, Jean join Angel and Stefano at the telescope.

16:00-21:00; Angel: no observations. Frozen rain and wind. Deicing working since yesterday night.

21:00-03:00; Stefano: wind and ice. The telescope is parked.


Jan. 18th, Wednesday

03:00-09:00; Angelos: no observations. Frozen rain went to snow before dawn; mild wind.

09:00-15:00, Sofia: no observations.

Let's take a little walk outside...


15:00-21:00 Frederic: no observations.

21:00-03:00 Stefano: Wind and snow. The telescope is completely frozen and parked. Few minutes after 2 a.m. local time, the taumeter drops suddenly to tau=0.1. We take a look outside. The sky is completely clear, the fog and clouds are no more there, the powder snow does not "pollute" the air anymore, the snowcats are working on the mountain, the stars shine, the thermometer marks -14 Celsius. Gloria increases the power of the de-icing. While we wait to have the 30m operational again, it will soon be time to chase the comet with binoculars and camera.

Jan. 19th, Thursday

03:00-09:00; Angelos: no observations. Clear sky, 100% visibility, no wind after dawn. Pedro and Salvador inspected the telescope and optimised the deicing subsystems by increasing power to some sectors of the backside. The Sun is our best ally right now.

attachment:30m_20220119_1.jpeg attachment:30m_20220119_2.jpeg

09:00-15:00; Sofia: no observations

15:00-21:00; Frédéric: no observations.

21:00-03:00; Stefano: let it snow, let it snow, let it snow...

attachment:IMG_7172_mod_sm.JPG attachment:IMG_7176_mod_sm.JPG

An old dream come true. We start a new tradition: Thursday midnight homemade pesto!!! Thank you to Angel for supporting the idea, to the 30m staff for finding the basil and to all those who stayed up late and joined :-)

attachment:IMG_7221_sm.JPG attachment:IMG_7223_pesto_sm.JPG

Jan. 20th, Friday

03:00-09:00; Angelos: no observations. It keeps snowing.

09:00-15:00; Sofia: no observations. Sunny. Deicing in progress.

15:00-21:00; Frédéric: no observations. Deicing stopped at 19:05.

21:00-03:00; Stefano: [DAQ restarted at 19:45; bad KIDS: 50 (ar1), 7 (ar2), 69 (ar3)]; we start observing at 20:00 after 48 hours of de-icing. The telescope is miraculously on focus. Calib_scan on CRL618 (scan nr 5). First project 105-22, half repetition of B213-C6 (because then it's at too high elevation). Then I move to 126-22 but the scripts do not work. ~30 minutes lost trying to figure it out. The a set of calibration, including a beam map and a calib_1scan on Uranus (scans number 26 and 27). Then back on 105-22, that meanwhile has come down enough again and I perform 5 repetitions. The sky is then degrading and the shift ends.

Today's explorations:

attachment:IMG_7299_sm.JPG attachment:IMG_7326_ok2_mod_sm.JPG attachment:IMG_7329_ok2_mod_sm.JPG

Jan. 21st, Saturday

03:00-09:00; Angelos: Monitoring the stability of the opacity. It was varying significantly, with fluctuations of +/- 10 Jy/b (2 mm). We shifted to EMIR and tried to calibrate with the receiver but we encountered a small setback with the carousel/dichroics. Cold load in EMIR calibrations was not possible for EMIR. We switched back to NIKA2. DAQ started at 05:15; bad KIDS: 62 (ar1), 7 (ar2), 67 (ar3). Again, we tried to monitor the stability of the opacity, but the conditions remained unstable. The operator performed more tests in order to fix the issues with the EMIR calibrations. We keep filling time by monitoring the stability. Instabilities increased above 50Jy/b. Humidity reached 100%. Dense fog. We stopped the telescope at 07:00.

09:00-15:00; Sofia: Observations started at 11:00 because of bad opacity conditions. We started observing with EMIR on project 124-22, completing the first two sources scans with good pointing. At 12:00 we continued with the other two scans, but at 12:45 we had to stop because of the increase of the opacity (0.5 at 13:00). Resuming observations at 14:20, pointing is now good enough, so we complete project 124-22 by running the last two scans. Last scan launched at 14:50 (lasts 40 mins)

15:00-21:00; Frederic & Jean: The weather is getting bad. We canceled the last scan at 3:02pm. We resumed EMIR observations at 6:30pm. We observed project 125-22. We performed four scans (Tile 5, beta 5 to 8). We swicthed to NIKA2 at 8:00pm.

tau225=0.17 @ 17:30. tau is stable since 1.5h and prediction is good -- EMIR 109 GHz observations. Are there reasons not to switch to NIKA2 ? <CK>

Several episodes of dense fog and clouds completely embedding the telescope. Before switching to NIKA2, we waited for stabilization of the conditions, so to avoid having to wait with NIKA2 on the telescope inactive if such episodes occur again [since switching back and forth takes a significant amount of time]. <SB> tau225 indicated that clouds are only local and not thick at all. But probably the taumeter does its skydips at the sweet spot of the atmosphere. Good luck !

@20 Local time: switched back to NIKA2; DAQ restarted, bad KIDS: 46 (ar1), 7 (ar2), 63 (ar3). After the initial setup, we start observing 126-22. The new scripts work fine.

21:00-03:00; Stefano: I take over and observe one repetition of the scripts nr. 1, 2 of the project 126-22. Then 159-22, 10 repetitions on NGC1569. Possible midnight (local time?) problem in scan 20230121s141. Beammap scan 20230121s139. Then I switch to 199-16 and observe 3 repetitions of PLCKG227. Finally I switch to 192-16 and start observing COSMOS right before the shift is over.

Midnight Fusion:

attachment:IMG_9518_mod_sm.JPG attachment:IMG_9500_mod_sm.JPG

Jan. 22nd, Sunday

03:00-09:00; Angelos: At the start of my shift I took over 192-16 and did one repetition of COSMOS. Then I continued with 160-16, 6 repetitions on NGC4254 which completed the IMEGIN program. Finally, I changed to 169-22 and started observing GJ526North. I did two repetitions until the end of my shift.

09:00-15:00: Sofia: DAQ restarted, bad KIDS: 48 (ar1), 6 (ar2), 66 (ar3). Skydip at 9:00, pointing on source 1226+023, no need for focusing. Opacity is 0.12 at 9:40, except for a spike of 0.3 at 9:00. Observing cluster PSZ2G094, currently 0/6 repetitions -> plan to do 6/6 (~2 hrs). Launching second set of scans of cluster PSZ2G094 at 10:40. Pointing (on source 1418+546) is good, no need for focusing. At 11:04 we start a new pointing on source 1633+382 and perform a focus. Scans on cluster 94 interrupted at 4th. Scans 5 and 6 completed at 12:10. Then pointing on 1637+574 and focus. Targeting source PSZ2G081, running repetitions 6-7-8. Then switch to project 171-22 and targeting cluster PSZ2G085. Opacity is great: 0.09 at 12:45. Scans of cluster PSZ2G085 started at 14:00, running repetitions 5-6-7.

15:00-21:00: Frederic & Jean: continuing project 171-22 on PSZ2G085. We started with a pointing/focus/pointing on 1823+568. Everything is looking good, except the beam is a bit large (18.1x18.9"). After two repetitions on source, we check the pointing on a brighter calibrator (1807+698). At 16:40, we switched to nikaw-22, performed a skydip and pointing/focus on MWC349 +calib1scan on MWC349, NGC7027, CRL6822 and Neptune. We resumed project 171-22 at 17:36, on source ACTJ0022 (3 repetitions). Overall, conditions were good (tau<0.1 & peak-to-peak < +/- 2 Jy/beam). At 6:42pm, we performed a pointing/focus/calib_1scan on CRL618. Focus changed a lot, from -1.35 to -0.7 (we performed 2 in a row). We then moved to GASTON's B213E_N at 19:15. We did 1 iteration, at which point B213E_N was too high. We started observing project 159-22 at 20:10.

21:00-03:00 Stefano: I take over, continue 159-22 on NGC 1569 until completion and then switch to calibrations: Uranus, beam map scan 251, calib_1scan 252. Project 126-22 follows, with one repetition for the maps nr. 3, 4, 5. I start then to observe COSMOS (project 192-16), but we have to stop and park telescope after few minutes because the wind raises suddenly above ~20 m/s. And it seems to be destined to increase further. The taumeter says 0.09. SIGH.

Chasing a young Moon:


Jan. 23rd, Monday

03:00-09:00; Angelos: no observations. High wind.

09:00-15:00: Sofia: no observations for the moment because of strong wind. At 12:30 wind speed dropped at ~8 m/s. DAQ restarted, bad KIDS: 45 (ar1), 6 (ar2), 49 (ar3). Observations starting at 12:45, pointing on source 1641+399. Focus, pointing again, then switch to project 199-16 and observe cluster 81 (3 scans left to do). Project 199, cluster 81 completed. Running last two scans of project 171, cluster 85.

15:00-21:00: Frederic: waiting for the last scan on 171-22/PSZG085 to finish. Low opacity (tau~0.1), but large fluctuations (rms~7-8 Jy/b). We moved to 171-22/ACTJ0022 at 16:00 (6 scans=>finished), we took the opportunity to do a calib_1scan on Neptune, passing by. Series of calibrations on CRL2688, NGC7027 and CRL618 at 18:25. We started to observe GASTON/B213E_N2 (1 repetition) at 19:10. Skydip at 20:00, followed by a calibration on ALFORI. We started a repetition on Orion B (map number 6) at 20:55.

21:00-03:00 Stefano: I take over and continue on 126-22 and perform 1 repetition on maps 6, 1, 2. Then calibrations: calib_1_scan (s114) and beam map (s122). GASTON follows, with one repetition on B213E_N2. The 199-16 with 2 repetitions on PLCKG227 (cluster completed!). Then 192-16, observing COSMOS (2 repetitions). Note that the pointing offsets in elevation seem to be drifting. During the first half of the night they reached 8.5 arcsec and the 11 arcsec.

Bye, all. It's always difficult to leave. I really feel at home here. See you next time, I hope.
One last treat before leaving Pico Veleta <:(
The moon dancing with Venus and Saturn

Jan. 24th, Tuesday

03:00-09:00; Angelos: started my shift with 2 repetitions of COSMOS. During the observations of both repetitions the opacity increased suddenly and dropped back to its prior value. Overall in both repetitions the opacity was quite stable around 0.21. After a skydip, I started observing PSZ2G126 of 171-22. I did 3 repetitions of that particular target. Last [almost] observation of week3 NGC4214 of 159-22 (5 repetitions).

Thank you all for your support and wonderful time spent in IRAM. Hope we meet again :)

09:00-10:30; Angel: Opacity about 0.12 at the end of the week. Getting pointing/focus on 1502+106. Quite stable: p2p variations ~ +/- 0.5 Jy/b (2 mm). Observing GJ526North for project 169-22: 1 repetition + pointing at the end. Finishing with a calib_1scan on MWC349 (s88) and NGC7027 (s89) + skydip (s90).

Thanks so much to those who have made possible the development of this pool week.

DailyReportsNika2Pool170123 (last edited 2023-01-24 20:13:18 by NikaBolometer)